OpenStudy (anonymous):

x2-2x-13=0 breaking this down following the 6 steps I end with the right side =x=1 and left side x=-4 is this right?

OpenStudy (anonymous):

You have to use the Quadratic Formula to solve this because the solutions will be imaginary.

OpenStudy (anonymous):

thats the 6 steps im talking about just looking for confirmation on the answer cause my teach is a harda**. Thanks

OpenStudy (anonymous):

ok ill post what i get.. just a sec

OpenStudy (anonymous):

Wait a minute

OpenStudy (anonymous):

Thats not right

OpenStudy (anonymous):

\[\frac{2 \pm \sqrt{(-2)^2 - 4(1)(-13)}}{2}\] \[\frac{2 \pm \sqrt{56}}{2}\] \[\frac{2 \pm 2 \sqrt{14}}{2}\] \[1 \pm \sqrt{14}\] \[x = 1 + \sqrt{14}\] \[x = 1 - \sqrt{14}\]

OpenStudy (anonymous):

There i finally got it right >.<

OpenStudy (anonymous):

And this site confirms that under the section "Roots" http://www.wolframalpha.com/input/?i=x^2-2x-13

OpenStudy (anonymous):

ill check out the site you posted thanks for the answers and help though she doesnt want it in the simple form i tell you she's a harda**? lol thanks again